OpenStudy (anonymous):

Digit manipulation proof, elucidate me. \[0.999... = 1\] I understand how the infinite series and sequences proof works however, why is the digit manipulation proof valid? \[x = 0.999... \\ 10x = 9.999... \\ 10x = 9 + 0.999... \\ 10x = 9 + x \\ 9x = 9 \\ x = 1 \] I will expand in the thread below.

OpenStudy (anonymous):

We already know that \(x = 0.999...\). Step 2 where \(10x = 9.999...\) should not come after: \(10 + 0.999... = 9.999...\) Not what ever they did?

OpenStudy (anonymous):

As in we know what \(x\) is, what are we doing by manipulating it?

OpenStudy (zzr0ck3r):

10*0.99999.....= 9.9999999.....

OpenStudy (zzr0ck3r):

10+.99999= 10.999999

OpenStudy (anonymous):

Oops silly my mistake.

OpenStudy (zzr0ck3r):

:)

OpenStudy (anonymous):

However why are we solving for \(x\) since we already know what the value is?

OpenStudy (zzr0ck3r):

where did you read this proof?

OpenStudy (anonymous):

@zzr0ck3r here: https://en.wikipedia.org/wiki/0.999...#Digit_manipulation

OpenStudy (zzr0ck3r):

tell them to prove that 10*.9999.... = 9.99999....

ganeshie8 (ganeshie8):

thats a standard trick for converting repeated decimals to fractions

OpenStudy (anonymous):

@ganeshie8 care to share this standard trick?

ganeshie8 (ganeshie8):

if you want more practice, try converting \(0.123123123123123\ldots\) to fraction follow the exact same steps :)

OpenStudy (mathmath333):

http://www.mathsisfun.com/9recurring.html

OpenStudy (anonymous):

However this step: \(9x = 9\) We know that \(x = 0.999...\) \[9x = 9 \\ 9 * 0.999... \ne 9\] It should return a value less than 9 right? For example: \[9 * 0.99 = 8.91\]

ganeshie8 (ganeshie8):

9x = 8.999... = 9

OpenStudy (shadowlegendx):

Round up dat figure ;)

ganeshie8 (ganeshie8):

\(0.99999999999999999999 \ne 1\) however \(0.999\ldots = 1\)

OpenStudy (anonymous):

I know that to be true via the infinite series and sequences proof. However how is the current proof true?

OpenStudy (shadowlegendx):

Orange = Black

ganeshie8 (ganeshie8):

after reaching 9x = 9 you move ahead, dont look back. divide 9 both sides

ganeshie8 (ganeshie8):

if you look back, you will get a repeating decimal again and you will be in circles

OpenStudy (anonymous):

\[x = 0.999... \\ 10x = 9.999... \\ 10x = 9 + 0.999... \leftarrow here\\ 10(0.999) \ne 9 + 0.999... \\ 10x = 9 + x \\ 9x = 9 \\ x = 1\] Never mind that, so the proof can only be read forward not backward? Fascinating.

ganeshie8 (ganeshie8):

you read it back wards if you want to prove that 1 = 0.999999....

OpenStudy (shadowlegendx):

Prove that an untrue statement is true o-o

ganeshie8 (ganeshie8):

there is a mistake here : \[ x = 0.999... \\ 10x = 9.999... \\ 10x = 9 + 0.999... \leftarrow here\\ \color{red}{10(0.999) \ne 9 + 0.999... \\} 10x = 9 + x \\ 9x = 9 \\ x = 1 \]

ganeshie8 (ganeshie8):

both sides will be equal if you replace "0.999" by "0.999..." on left hand side

OpenStudy (anonymous):

\[1:x = 0.999... \\ 2:10x = 9.999... \\ 3:10x = 9 + 0.999... \leftarrow here\\ 4:10(0.999) \ne 9 + 0.999... \\ \text{Step 4 makes no sense since we know from the beginning that x = 0.999...} \\ 5:10x = 9 + x \\ 6:9x = 9 \\ 7:x = 1\] Oh wait I see my mistake, I jumped, I took from step 1 when I was working backwards instead of taking from step 7. Thank you @ganeshie8 it makes sense now.

ganeshie8 (ganeshie8):

looks good!

ganeshie8 (ganeshie8):

food for thought : ...999999999999999999999 +1 -------------------------- ...000000000000000000000

ganeshie8 (ganeshie8):

it may not be that exciting to you as you're good with computers and programming

OpenStudy (anonymous):

Is repeating numbers any different than the concept of infinity?

ganeshie8 (ganeshie8):

i think it can be explained easily using overflow/carry...

ganeshie8 (ganeshie8):

perform below operation on a 4 bit computer and see the contents of accumulator : ``` ADD F 1 ```